OpenStudy (anonymous):

The perimeter of a sector is 20 units. If the length of the intercepted arc is 10 units, find the area of the sector.

sam (.sam.):



sam (.sam.):

Perimeter=r+r+s=2r+s

OpenStudy (anonymous):

or P = r theta + 2 r

sam (.sam.):

Yes that as well, from s=rθ \[20=2r+\theta r...(1)\] Area, \[A=\frac{1}{2}r^2 \theta...(2)\]

sam (.sam.):

\[P=2r+\theta r~~~~~s=r \theta \\ \\ 20=r(2+\theta)~~~10=r \theta \\ \\ r=\frac{20}{2+ \theta}\] ------------------------------ \[10=\frac{20 \theta}{2+\theta} \\ \\ \theta=2\] \[r=5\] ------------------------------ From area, \[A=\frac{1}{2}r^2 \theta \\ \\ \]

OpenStudy (anonymous):

Thank you so much! I was a bit confused in the problem awhile ago.

sam (.sam.):

yw :)